four years ago, a man's age was three years more than thrice his son's age. fifteen years hence, his age will be five years less than twice his son's age . find the man's present age.
Answer by stanbon(75887)

four years ago, a man's age was three years more than thrice his son's age. fifteen years hence, his age will be five years less than twice his son's age . find the man's present age.
---
Equations:
m-4 = 3(s-4)+3
m+15 = 2(s+15)-5
-----
Rearrange:
m - 7 = 3s - 12
m +20 = 2s + 30
------------------
Modify:
m - 3s = -5
m - 2s = 10
---
Subtract and solve for "m":
son's age = 15
---
Solve for "m":
m - 2s = 10
m -2*15 = 10
man's age = 40
=====================
Cheers,
Stan H.
======================